OpenStudy (anonymous):

how do I solve 2a^2-a-21=0 by factoring?

OpenStudy (anonymous):

You have to first observe the -21. You have to find numbers which you can multiply together to get 21. Those numbers are: 1 * 21 3 * 17 Now observe the the -a and the 2a^2. You know you need a factor that looks like (2a + b)(a + n) What combination of numbers will give you a -a as your middle term? I know you're familiar with factoring equations like x^2 + 3x + 2 = (x+2)(x+1) = x^2 + x + 2x + 2 Same thing applies here. You just need to choose between these following options. Only one of them is right but you'll find that out by FOILing. I know the answer but you should FOIL it yourself to find the right answer. I hope that helps. (2a^2 - 21)(a - 1) (2a^2 - 1)(a - 21) (2a^2 - 7)(a + 3) (2a^2 - 3)(a + 7)

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

You're welcome!